In a wallet are one-peso and five-peso coins. The number of coins sum up to 22 but its value sums up to 50. How many one-peso and five-peso coins are in it?
Answer by ikleyn(52781)

.

F = # of five-peso coins

22-F = # of one-peso coins


Total money equation


    5F + (22-F) = 50

      4F        = 50-22 = 28

       F                = 28/4 = 7.


ANSWER.  7 five-peso coins, and the rest  (22-7) = 15 are one-peso coins.

Solved.